OpenStudy (anonymous):

I'm doing a bit of maths revision before a university degree. Stuck on an algerbra question. Rearrange the following equation in terms of y: 2x + 3y = 4(x + y) y = How do I work this out to get y please?

OpenStudy (dumbcow):

step 1: distribute right side 2x + 3y = 4x +4y step 2: combine like terms 2x + 3y = 4x + 4y -4x -3y -4x -3y ------------------------ -2x = y step 3: divide by coefficient of y y/1 = -2x/1 y = -2x

OpenStudy (dumbcow):

Note: step 3 is really only needed when coefficient is not 1